OpenStudy (tainted):

Solve and Show Work) 1. √8x+1=5 2. √x-7+5=11 3. -4√x+9=20 4. √6x+5=√53 5. √x+2+4=x

OpenStudy (anonymous):

it's not difficult ! i will hepl you : try to work it is better for you too : so for the first one, you have V8x+1=5 so V8x=5-1 V8x=4 x=...

OpenStudy (anonymous):

Do you need help?

OpenStudy (anonymous):

sqrt(8x)+1=5 Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides. sqrt(8x)=-1+5 Add 5 to -1 to get 4. sqrt(8x)=4 To remove the radical on the left-hand side of the equation, square both sides of the equation. (sqrt(8x))^(2)=(4)^(2) Simplify the left-hand side of the equation. 8x=(4)^(2) Expand the exponent (2) to the expression. 8x=(4^(2)) Squaring a number is the same as multiplying the number by itself (4*4). In this case, 4 squared is 16. 8x=(16) Remove the parentheses around the expression 16. 8x=16 Divide each term in the equation by 8. (8x)/(8)=(16)/(8) Simplify the left-hand side of the equation by canceling the common factors. x=(16)/(8) Simplify the right-hand side of the equation by simplifying each term. So x=2

OpenStudy (anonymous):

This is for number 2 sqrt(x)-7+5=11 Add 5 to -7 to get -2. sqrt(x)-2=11 Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides. sqrt(x)=2+11 Add 11 to 2 to get 13. sqrt(x)=13 To remove the radical on the left-hand side of the equation, square both sides of the equation. (sqrt(x))^(2)=(13)^(2) Simplify the left-hand side of the equation. x=(13)^(2) Expand the exponent (2) to the expression. x=(13^(2)) Squaring a number is the same as multiplying the number by itself (13*13). In this case, 13 squared is 169. x=(169) Remove the parentheses around the expression 169. So, x=169

OpenStudy (anonymous):

I can't really tell if you wrote your equations correctly because they don't have parentheses in them.

OpenStudy (anonymous):

is the first equatin √8x+1=5 or √(8x+1)=5?

OpenStudy (anonymous):

*equation.

OpenStudy (tainted):

They do not have any parentheses in them

OpenStudy (anonymous):

You are wrong.

OpenStudy (anonymous):

For the first one, it is sqrt (8x+1)=5 second one: sqrt(x-7)+5=11 If you don't put that in your question, you'll get completely different answers.

OpenStudy (tainted):

Oh I was not aware of that.

OpenStudy (anonymous):

If the square root extends over those numbers, you either show that by drawing a longer square root, or put the part of the equation in parentheses. Just a future heads up. BTW, then the answers that I gave you are incorrect.

OpenStudy (anonymous):

Since the equation that I solved was different than the one in the picture.

OpenStudy (anonymous):

Solution 1. sqrt(8x+1)=5 To remove the radical on the left-hand side of the equation, square both sides of the equation. (sqrt(8x+1))^(2)=(5)^(2) Simplify the left-hand side of the equation. 8x+1=(5)^(2) Expand the exponent (2) to the expression. 8x+1=(5^(2)) Squaring a number is the same as multiplying the number by itself (5*5). In this case, 5 squared is 25. 8x+1=(25) Remove the parentheses around the expression 25. 8x+1=25 Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides. 8x=-1+25 Add 25 to -1 to get 24. 8x=24 Divide each term in the equation by 8. (8x)/(8)=(24)/(8) Simplify the left-hand side of the equation by canceling the common factors. x=(24)/(8) Simplify the right-hand side of the equation by simplifying each term. So, x=3